OpenStudy (anonymous):

Rationalize denominator and simplify. A. 2/ sqrt 7+ sqrt 5 B. 6/sqrt 27

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{7}+\sqrt{5} } \times \frac{ \sqrt{7}-\sqrt{5} }{ \sqrt{7}-\sqrt{5} }\] this is what it means, notice the change in sign.

OpenStudy (anonymous):

Yes can u help me finish it

OpenStudy (anonymous):

Sure, so what did you get for the first one?

OpenStudy (anonymous):

Make sure you use parenthesis so you don't make mistakes

OpenStudy (anonymous):

Is the answer going to be sqrt 7sqrt5/12

OpenStudy (anonymous):

How did you get that?